OpenStudy (nsh4267):

Write the expression in the form a+bi. i(3+4i)^2

OpenStudy (anonymous):

expand (3+4i)^2 using the identity\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

then multiply it by i

OpenStudy (klimenkov):

Use \((a+b)^2=a^2+2ab+b^2\) and \(i^2=-1\).

OpenStudy (nsh4267):

so it should be 9+12i+16i^2 right?

OpenStudy (phi):

remeber i^2 is real = -1, so you can simplify some more

OpenStudy (phi):

remember i^2 is real = -1, so you can simplify some more however, you lost the bubble some where because your expression 9+12i+16i^2 = -7+12i does not match i*(3+4i)^2